OpenStudy (aromire):

An object starts from rest with an acceleration of 2.0 m/s2 for 3.0 seconds. It then reduces its acceleration to 1.0 m/s2 for 5.0 additional seconds. The total distance covered is

OpenStudy (ybarrap):

Using these (SUVAT) equations http://en.wikipedia.org/wiki/Equations_of_motion#Constant_linear_acceleration:_collinear_vectors We find for the 1st 3 seconds, it travels, $$ r_{3 sec}=r_0+v_0t+{at^2\over2}\\ r={2\times3^2\over2}=9~meters $$ Its velocity at the end of 3 seconds is $$ v=at=3\times3=9~m/s $$ So after 5 seconds, it has traveled (using first equation above) $$ r_{5 sec}=9+3(5)+{1(5)^2\over 2}~meters $$

OpenStudy (aromire):

Thank you ybarrap but in the options, the possible answers are 53 meters, 38 meters, 11 meters, 9 meters or 7.6 meters. None of it match your own answer please I need help. Thank you very much!

OpenStudy (aromire):

52 meters I mean.

OpenStudy (anonymous):

for what ybarrap has said using the equation V=U+at the final velocity at the end of 3 second is 6m/s2 but not 9

OpenStudy (anonymous):

for the distance s2 s=ut +1/2 at^2 s=6(5) +1/2 (-1)5*5 s=`17.5 adding both we get 26.5

OpenStudy (shamim):

For 3 sec use S=ut+0.5at^2 S=0*3+0.5*2*3^2=?

OpenStudy (shamim):

Now after 3sec the final velocity will b v=u+at v=0+2*3 v=6m/s Right?

OpenStudy (shamim):

Now for additional 5 sec S=ut+0.5at^2 S=6*5+0.5*1*5^2 S=?

OpenStudy (shamim):

Now add those 2 s. U will get total distance travelled

OpenStudy (shamim):

Response plz!!!!!!

OpenStudy (aromire):

Hi shamim, the answer for S for the 3 sec is 9 m and the S for the additional 5 sec is 42.5 and after adding 9 and 42.5 I got 51.5 which is approximately 52. And that's part of the option that was given in the question. Thank you very much.